OpenStudy (anonymous):

What is the length of the missing side of the right triangle shown below? square root of 74 Two times the square root of 15 10 2

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Where is the pic?

OpenStudy (anonymous):

use pythagoras theorem

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

I believe its the first choice. square root of 74

OpenStudy (anonymous):

Let the hypotneus be x x = \[x = \sqrt{5^2 + 7^2}\]

OpenStudy (anonymous):

yahoo is that right square root of 74

OpenStudy (anonymous):

@pumpgirls thats it

OpenStudy (anonymous):

\[= \sqrt{25 + 49} = \sqrt{74}\]

OpenStudy (anonymous):

yes...@pumpgirls

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Welcome